Use inductive reasoning to determine the next three numbers in the pattern: 0,1,3,6,10,15...
Answer by stanbon(75887)

Use inductive reasoning to determine the next three numbers in the pattern: 0,1,3,6,10,15...
----
0+1 = 1
1+2 = 3
3+3 = 6
...
...
15+6= 21
21+7= 28
28+8= 36
=====================
Cheers,
Stan H.